OpenStudy (anonymous):

Part 1: Create your own term using a rational exponent. Part 2: Convert it to a radical expression. Part 3: Explain, in complete sentences, how the expression was converted

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@waterineyes @nbouscal

OpenStudy (anonymous):

See, let me make the term as : \[\large x^y\] Okay??

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Or you want the rational exponent so you can take like this also: \[\large x^{\frac{1}{y}}\] Is it okay??

OpenStudy (anonymous):

Sure no problem. Let's use x^1/y

OpenStudy (anonymous):

So, in radical form it is converted as: \[\huge x^{\frac{1}{y}} = \sqrt[y]{x^1}\] Getting it or not??

OpenStudy (anonymous):

Yes I get it. Okay.

OpenStudy (anonymous):

I just don't understand how I would explain how I converted it.

OpenStudy (anonymous):

See, Numerator becomes the exponent and denominator becomes the nth root... More examples: For: \[\huge x^{\frac{2}{3}} = \sqrt[3]{x^2}\]

OpenStudy (anonymous):

K tyvm :)

OpenStudy (anonymous):

I give you the general formula for this: \[\huge \color{blue}{x^{\frac{y}{n}}= \sqrt[n]{x^y}}\]

OpenStudy (anonymous):

Welcome dear.. All you have to do is to replace y and n according to the question..

OpenStudy (anonymous):

K. Tyvm :) I just didn't know how to explain it (Part 3) but now I do :D